# Persistent Backdoor Attacks in Continual Learning

### v1

@@ 2.2 Backdoor Attacks
A backdoor attack is an adversarial technique where an attacker injects malicious patterns, known as triggers, into a small subset of the training data, assigning these data points incorrect labels [4, 42, 49, 50]. During the inference, the backdoored model produces incorrect outputs whenever the specific trigger is present in the input. This allows the adversary to control the model’s decisions on triggered inputs, while keeping the model’s general performance intact. Backdoors are a more targeted form of attack compared to universal adversarial perturbations [31, 25, 29, 57, 15]. While both cause a model to misclassify inputs to an attacker-chosen label, backdoors require modifying both the model and the input, but offers more flexibility. Moreover, backdoor attacks differ from data poisoning, where the attacker’s primary goal is to degrade the overall model performance.
@@ 3 Observation
Insight 1 – Algorithmic-level Analysis: At the highest granularity level, we analyzed various CL algorithms to characterize their overall behavior and parameters’ variation as they learn new tasks. Figure 3 illustrates the parameter variations for six CL algorithms, calculated using the aforementioned approach. From this figure, it is evident that regularization-based algorithms exhibit a lower degree of variation in their parameters compared to replay-based algorithms. The rationale behind this phenomenon is that regularization-based algorithms, such as SI, EWC, XdG, and LwF, tend to penalize drastic changes to the parameters learned in prior tasks, imposing a high cost for modifications to the learned model. In contrast, replay-based algorithms, such as DGR and A-GEM, do not impose such constraints. Instead, these algorithms retain the knowledge of the previously learned tasks by replaying and learning from prior data distributions without penalizing parameter updates. As a result, the model experiences a higher degree of variation in its parameters’ values. For instance, our quantification revealed that the average variation in the parameter values of the regularization-based EWC algorithm is only 3.93, while the replay-based A-GEM algorithm shows an average value of 9.52.
@@
In neural networks, as the depth increases, layers transition from learning low-level features like edges and textures to high-level, abstract representations such as shapes and objects. As a result, in a CL context, it is expected that the parameters of the earlier layers remain relatively stable when learning new data distributions, while those in the final layers exhibit more significant changes–as we observed in the EWC algorithm. Contrary to this understanding, five other algorithms demonstrated the unexpected behavior of higher variation in the earlier layers. We attribute this to the possibility that the features learned in the earlier layers are more sensitive to the new task data distribution, especially if the new tasks require fundamentally different low-level features. This sensitivity can lead to greater variation as the network adjusts to accommodate these new features.
Figure 4: Our layer-wise analysis revealed that for the EWC algorithm, the last layers experience larger variations compared to the previous two layers, whereas for the LwF and A-GEM algorithms, the last layers exhibit smaller variations than the preceding layers.
@@
Figure 5: Parameter variation of the 49 selected neurons from the model’s third layer (using LwF algorithm). $T_{12}$ shows parameter values’ variation between Task 1 and Task 2, $T_{13}$ between Task 1 and Task 3, $T_{14}$ between Task 1 and Task 4, and $T_{15}$ between Task 1 and Task 5. A subset of neurons, marked with the red icon, demonstrate high stability as the model evolves across all tasks.
@@
Insight 3 – Neuron-level Analysis: Finally, we analyzed neuron stability as the model evolved, focusing on the third layer of the LwF algorithm due to its greater stability. Given the high dimension of parameters, we selected 49 out of 400 neurons for better illustration by dividing the parameter variations into seven ranges and sampling seven neurons from each range (Figure 5). We then quantified the variations in these neurons’ values between Task 1 and all subsequent tasks (e.g., $T_{13}$ represents the changes in neuron values from Task 1 to Task 3), and visualized the results in each sub-figure of Figure 5. This revealed a pattern: a small subset of neurons, marked with a red icon, remains remarkably stable with almost no change throughout the entire training process across all tasks. In contrast, most other neurons exhibit varying degrees of change, with some maintaining relative stability between two or three consecutive tasks before undergoing a significant shift in values.
@@ 4.1 Threat Model and Assumptions
The rise in deep learning’s popularity has turned platforms like GitHub and Hugging Face into repositories for hundreds of thousands of open-source models. These models are publicly available either as pre-trained weights, ready for immediate use in various tasks, or as source code, which developers can adapt for specific applications. This accessibility has led to the widespread integration of open-source code, driven by the need to reduce costs and accelerate the development of large models [1]. For instance, Hugging Face reported over 600 million model downloads in August 2023 alone [6]. However, the widespread use of source code from diverse repositories has also increased the risk of security vulnerabilities, as highlighted by recent high-profile attacks that have exposed the dangers of compromised code [38, 3]. Given these threats, we consider two specific attacks–Blind Task Backdoor and Latent Task Backdoor–in which the adversary targets backdooring a specific task(s) in the model. These threats are particularly stealthy, as the backdoor remains dormant until the targeted task is executed, significantly reducing the risk of their detection.
@@
The blind task backdoor (BTB), which we will refer to as the blind backdoor for brevity, is a type of code poisoning attack where the adversary has no access to the training data and cannot control or observe the model training process. Similar to [3], the adversary maliciously manipulates the machine learning codebase, referred to as the attack code, specifically targeting the calculation of the loss value during training. The attack code also includes a module for generating poisoned data samples by injecting triggers into a small subset of the training data and assigning incorrect labels, as detailed in [3]. As a result, the adversary remains blind to both the training process and the data. In continual learning, the same attack code is used across all tasks, allowing the backdoor to be injected at any point in the learning process without the adversary directly interacting with the model during training. This attack is particularly effective when ML code sourced from a repository is used unchanged to train a model, allowing the adversary to embed malicious behaviors without the developer’s knowledge, thereby compromising the application’s integrity. We note that loss computation in deep learning models is context-dependent, varying with architecture, data modality, tasks, and the CL algorithm. This diversity often results in a single codebase employing multiple loss functions, making it challenging to detect malicious alterations through code review alone, as detailed in [3].
Unlike the blind backdoor, which does not involve access to the training process, the latent task backdoor (LTB), referred to as the latent backdoor, is a sophisticated attack methodology where the adversary can manipulate the training process of only one task. Specifically, the adversary modifies the loss computation of only the target task and minimally adjusts a subset of the model’s parameters during that task’s training. As a result, the training remains benign for all other tasks, with no code poisoning or loss manipulation in either prior or subsequent tasks. This attack is particularly pertinent when developers directly use pre-trained weights to build an application, fine-tune a custom model, or source pre-trained models as a foundation for learning additional tasks or transfer learning [52]. We note that detecting changes in pre-trained weights is challenging due to the randomness introduced by factors like the optimizer, initial seed, and data variations. In Section 6, we present empirical results demonstrating how these attacks evade backdoor detection mechanisms.
@@
In summary, for the blind backdoor, we assume the adversary compromises only the loss calculation in the ML code without controlling the model training. While the adversary may have some knowledge of the model architecture and general data domain, they have no access to the training data and no knowledge of the hyperparameters. In contrast, for the latent backdoor, we assume the adversary has control over the training process of a single task.
@@ 5.1 Blind Task Backdoor
As mentioned in Section 4.1, in Blind Task Backdoor (i.e., blind backdoor), the adversary generates the attack code by manipulating the loss value computation of the ML code, indirectly impacting the training process of the model across all tasks. For this attack to be successful, the modification of the loss computation should achieve a high attack success rate on the triggered inputs on the current and future tasks, and maintain high classification accuracy on the clean inputs across all tasks.
To achieve this, we devise a customized blind loss, aiming to effectively solve this multi-objective optimization problem, as detailed in Eqn. 1:
@@ 5     Compute $F\leftarrow\nabla_{\theta}^{2}\ell$ for each $\theta$
6 $\triangleright$ Select Stable Neurons from the overall $n$ neurons
@@ 11     $\ell_{latent}\leftarrow L_{i}(\theta(x^{+}),y^{+})+\text{ReLU}(L_{i}(\theta(x),y)-\epsilon)$
$\underset{\theta,\lambda,\mu}{\min}\,(\ell_{blind}^{i}+\sum_{j}\lambda_{j}\delta+\frac{\mu}{2}\sum_{j}\delta^{2})$ (3)
@@ 5.2 Latent Task Backdoor
DFM analysis enables us to evaluate neurons’ contributions to the classification performance of the task controlled by the adversary by assessing the loss function’s sensitivity to variations in neurons. Interestingly, our in-depth analysis revealed a significant overlap between the stable neurons and those identified by DFM analysis as crucial for the task’s classification performance. As a result, we select the top important neurons based on the importance matrix $F$, forming the candidate set, with $\kappa$ approximately corresponding to the 98th percentile of $F$ in our experiments. To maintain the attack’s stealth and persistence, we select a proportion $P\%$ of the candidate set for trigger embedding (line 5). Our algorithmic-level observations in Section 3 revealed that replay-based algorithms experience more significant parameter changes than regularization-based algorithms, necessitating a higher percentage of neurons for trigger embedding in replay-based algorithms. In our experiments, we set $P$ to 70% of the candidate set for regularization-based algorithms and 90% for replay-based algorithms.
@@ 6 Experiments
To validate the effectiveness of our proposed blind backdoor (BTB) and latent backdoor (LTB) attacks, we conducted extensive experiments (details are in Table 4 in the Appendix) to evaluate: (i) static backdoor against six CL algorithms, (ii) dynamic backdoor (iii) physical backdoor, (iv) backdoor on text classification, (v) comparative analysis against other backdoor methods, (vi) backdooring different tasks in CL, and (vii) evading existing backdoor detection defenses. We perform backdoor attacks on task-based CL with the code developed based on [46]. The backdoor is embedded in task 1 unless stated otherwise. All experiments are conducted using PyTorch and executed on an NVIDIA A100 GPU. For the BTB attack, we set the number of iterations ($n$) to 300, $\alpha$ to $0.001$, $\beta$ to $0.0001$, and $\tau$ to be $0.05\times\ell_{j}$, where $\ell_{j}$ is the loss value of original task $j$. For the LTB attack, we set $v^{tigger}$ to 0.5, $\epsilon$ to 0.1 of the backdoor loss ($0.1\times L(\theta(x^{+}),y^{+})$), $\kappa$ to 98 percentile of all neurons, and $P$ to 70% and 90% for regularization-based and replay-based algorithms.
@@
CNN SplitMNIST $0.95$$\pm 0.05$ $0.92$$\pm 0.02$ $0.93$$\pm 0.03$ $0.91$$\pm 0.02$ $0.94$$\pm 0.01$ $0.91$$\pm 0.03$ $0.91$$\pm 0.05$ $0.92$$\pm 0.02$ $0.92$$\pm 0.04$ $0.93$$\pm 0.03$ $0.91$$\pm 0.07$ $0.93$$\pm 0.03$
PermutedMNIST $0.93$$\pm 0.04$ $0.82$$\pm 0.09$ $0.85$$\pm 0.09$ $0.80$$\pm 0.05$ $0.90$$\pm 0.03$ $0.82$$\pm 0.01$ $0.92$$\pm 0.06$ $0.85$$\pm 0.03$ $0.91$$\pm 0.04$ 0.26 ±$\pm 0.02$ $0.89$$\pm 0.03$ $0.81$$\pm 0.04$
SplitCIFAR10 $0.91$$\pm 0.02$ $0.76$$\pm 0.01$ $0.91$$\pm 0.03$ $0.83$$\pm 0.04$ $0.88$$\pm 0.04$ $0.85$$\pm 0.01$ $0.84$$\pm 0.04$ $0.81$$\pm 0.03$ $0.94$$\pm 0.03$ 0.52 ±$\pm 0.02$ $0.89$$\pm 0.06$ $0.86$$\pm 0.03$
ResNet18 SplitMNIST $0.93$$\pm 0.01$ $0.96$$\pm 0.03$ $0.91$$\pm 0.01$ $0.92$$\pm 0.04$ $0.93$$\pm 0.03$ $0.91$$\pm 0.02$ $0.91$$\pm 0.01$ $0.98$$\pm 0.04$ $0.89$$\pm 0.03$ $0.91$$\pm 0.02$ $0.90$$\pm 0.03$ $0.93$$\pm 0.04$
PermutedMNIST $0.93$$\pm 0.01$ $0.86$$\pm 0.06$ $0.89$$\pm 0.03$ $0.88$$\pm 0.04$ $0.92$$\pm 0.01$ $0.79$$\pm 0.03$ $0.93$$\pm 0.06$ $0.91$$\pm 0.07$ $0.91$$\pm 0.01$ 0.35 ±$\pm 0.06$ $0.85$$\pm 0.03$ $0.85$$\pm 0.01$
SplitCIFAR10 $0.90$$\pm 0.03$ $0.80$$\pm 0.04$ $0.90$$\pm 0.03$ $0.82$$\pm 0.02$ $0.82$$\pm 0.02$ $0.78$$\pm 0.03$ $0.90$$\pm 0.01$ $0.83$$\pm 0.03$ $0.89$$\pm 0.02$ 0.63 ±$\pm 0.03$ $0.93$$\pm 0.01$ $0.88$$\pm 0.07$

### v3

@@ 2.2 Backdoor Attacks
A backdoor attack is an adversarial technique where an attacker injects malicious patterns, known as triggers, into a small subset of the training data, assigning these data points incorrect labels [43, 50, 51]. During the inference, the backdoored model produces incorrect outputs whenever the specific trigger is present in the input. This allows the adversary to control the model’s decisions on triggered inputs, while keeping the model’s general performance intact. Backdoors are a more targeted form of attack compared to universal adversarial perturbations [31, 25, 29, 59, 15]. While both cause a model to misclassify inputs to an attacker-chosen label, backdoors require modifying both the model and the input, but offers more flexibility. Moreover, backdoor attacks differ from data poisoning, where the attacker’s primary goal is to degrade the overall model performance.
@@ 3 Observation
Figure 4: The layer-wise analysis revealed that for EWC, the last layers experience larger variations compared to the previous two layers, whereas for LwF and A-GEM, the last layers exhibit smaller variations than the preceding layers.
Insight 1 – Algorithmic-level Analysis: At the highest granularity level, we analyzed various CL algorithms to characterize their overall behavior and parameters’ variation as they learn new tasks. Figure 3 illustrates the parameter variations for six CL algorithms, calculated using the aforementioned approach. From this figure, it is evident that regularization-based algorithms exhibit a lower degree of variation in their parameters compared to replay-based algorithms. The rationale behind this phenomenon is that regularization-based algorithms, such as SI, EWC, XdG, and LwF, tend to penalize drastic changes to the parameters learned in prior tasks, imposing a high cost for modifications to the learned model. In contrast, replay-based algorithms, such as DGR and A-GEM, do not impose such constraints. Instead, these algorithms retain the knowledge of the previously learned tasks by replaying and learning from prior data distributions without penalizing parameter updates. As a result, the model experiences a higher degree of variation in its parameters’ values. For instance, the average variation in the parameter values of the regularization-based EWC algorithm is only 3.93, while the replay-based A-GEM algorithm shows an average value of 9.52.
@@
Figure 5: Parameter variations of 49 selected neurons from the model’s third layer using LwF. $T_{12}$ represents the parameter variations between Task 1 and Task 2, $T_{13}$ between Task 1 and Task 3, $T_{14}$ between Task 1 and Task 4, and $T_{15}$ between Task 1 and Task 5. A subset of neurons, highlighted with red icons, exhibits high stability as the model progresses.
In neural networks, as the depth increases, layers transition from learning low-level features like edges and textures to high-level, abstract representations such as shapes and objects. As a result, in a CL context, it is expected that the parameters of the earlier layers remain relatively stable when learning new data distributions, while those in the final layers exhibit more significant changes–as we observed in the EWC algorithm. Contrary to this understanding, five other algorithms showed the unexpected behavior of higher variation in earlier layers. We attribute it to the possibility that features learned in the earlier layers are more sensitive to the new task data distribution, especially when new tasks require fundamentally different low-level features. This sensitivity leads to greater variation as the network adjusts to accommodate new features.
@@
Insight 3 – Neuron-level Analysis: We also analyzed the behavior of the neurons as the model evolved, focusing on the third layer of the LwF algorithm due to its greater stability. Given the high dimension of parameters, we selected 49 out of 400 neurons for better illustration by dividing the parameter variations into seven ranges and sampling seven neurons from each range (Figure 5). We then quantified the variations in these neurons’ values between Task 1 and all subsequent tasks (e.g., $T_{13}$ represents the changes in neuron values from Task 1 to Task 3), and visualized the results in each sub-figure of Figure 5. This revealed a pattern: a small subset of neurons, marked with a red icon, exhibits small variations in their values throughout across training all tasks. In contrast, most other neurons exhibit varying degrees of change, with some exhibiting a significant shift in values in later tasks.
Insight 4 – Component Stability Analysis: Finally, we analyzed neuron stability by examining weight variations during the continual learning process. Using the A-GEM algorithm, we trained a ResNet18 model on the SplitCIFAR10 dataset over five tasks. After Task 1, we identified 200 stable neurons, important for Task 1, using the diagonal Fisher matrix and randomly selected 200 others for comparison. We then trained the model on four additional tasks and quantified the weight variations of these neuron groups after each task relative to their baseline values established in Task 1, with $\delta_{1-i}$ representing the variation between Task 1 and Task $i$.
Figure 6: The kernel density estimation of parameter variations between stable and randomly selected neurons. The mean value of stable and random neurons are 0.05 and 0.36, respectively. Notably, the gap between the two distributions widens as the model progresses through additional tasks, indicating increasing differentiation in their respective stability.
We visualized the weight variation of these two neuron groups using the kernel density estimation as shown in Figure 6. It is evident that random neurons (orange distribution) exhibit progressively larger mean variations across tasks, whereas stable neurons (blue distribution) remain relatively consistent. Quantitative results further support this observation: stable neurons exhibit a mean variation of 0.05 with a standard deviation of 0.22, whereas randomly selected neurons show a significantly higher mean variation of 0.36 and a standard deviation of 0.44. Notably, the gap between the two distributions widens as the model learns more tasks, indicating that stable neurons effectively retain their weights, while random neurons experience increasingly significant variations (refer to Figure 20 in Appendix 10 for per-neuron value variation).
To further investigate the root cause of the observed neuron stability, we quantified the overlap among neurons utilized during the training of individual tasks using the normalized intersection-over-union (IoU) score and visualized these intersections as a heatmap in Figure 7. The analysis revealed that the largest intersection, excluding intra-task comparisons, is 6.4%, observed between Tasks 4 and 5, suggesting a small degree of overlap. However, the overall low IoU scores indicate that tasks primarily rely on distinct sets of neurons. These results demonstrate that different tasks in Cl rely on predominantly disjoint sets of neurons, and the neurons critical to a task remain stable in learning subsequent tasks.
Figure 7: Neuron overlap heatmap showing the percentage of neuron intersection across five tasks. Each cell represents the normalized intersection-over-union score between two tasks. Diagonal values of 100% in self-comparison highlight that distinct neurons are exclusively utilized for different tasks.
@@ 4.1 Threat Model and Assumptions
The rise in deep learning’s popularity has turned platforms like GitHub and Hugging Face into repositories for hundreds of thousands of open-source models. These models are publicly available either as pre-trained weights, ready for immediate use in various tasks, or as source code, which developers can adapt for specific applications. This accessibility has led to the widespread integration of open-source code, driven by the need to reduce costs and accelerate the development of large models [1]. For instance, Hugging Face reported over 600 million model downloads in August 2023 alone [5]. However, due to the widespread adoption of platforms like GitHub, Twitter, and Flicker in the ML community, the attack vector has expanded beyond poisoned data to include poisoned weights and carefully crafted code snippets [38, 3]. Given such threats, we consider two specific attacks–Blind Task Backdoor and Latent Task Backdoor–in which the adversary targets backdooring a specific task. These threats are stealthy, as the backdoor remains dormant until the targeted task is executed, significantly reducing the risk of their detection.
@@
The blind task backdoor (BTB), which we will refer to as the blind backdoor for brevity, is a type of code poisoning attack where the adversary indirectly impacts the model training process without access to the training data. Similar to [3], the adversary maliciously manipulates the machine learning codebase, referred to as the attack code, specifically targeting the calculation of the loss value during training. The attack code also includes a module for generating poisoned data samples by injecting triggers into a small subset of the training data and assigning incorrect labels, as detailed in [3]. As a result, the adversary remains blind to both the training process and the data. In continual learning, the same attack code is used across all tasks, allowing the backdoor to be injected at any point in the learning process without the adversary directly interacting with the model during training. This attack is particularly effective when ML code sourced from a repository is used unchanged to train a model, allowing the adversary to embed malicious behaviors without the developer’s knowledge, thereby compromising the application’s integrity. We note that loss computation in deep learning models is context-dependent, varying with architecture, data modality, tasks, and the CL algorithm. This diversity often results in a single codebase employing multiple loss functions, making it challenging to detect malicious alterations through code review alone, as detailed in [3].
Unlike the blind backdoor, which relies on indirect manipulation of the entire training process, the latent task backdoor (LTB), referred to as the latent backdoor, is a more sophisticated attack method in which the adversary manipulates the training process of only one task. The weight modifications introduced during this task are designed to backdoor future targeted tasks, even when their training processes remain benign. This restriction on the adversary showcases that, in CL, as in classical ML pipelines, publicly shared weights can contain dormant and hidden functionalities that may be exploited in later tasks. The adversary achieves this by modifying the loss computation for a single task, leading to a controlled update of specific weights during the training process. The restricted threat model of LTBs (i.e., single-task manipulation) highlights the risks developers face when directly using pre-trained weights to build applications, fine-tune custom models, few-shot learning [42], or perform transfer learning [54]. It also underscores that backdooring poses a significant threat in pipelines relying on untrusted foundation model weights for learning additional tasks [53], even when the training pipeline itself is verified and trusted.
We emphasize that in scenarios where the training pipeline is developed by legitimate entities and remains benign, the blind backdoor becomes inapplicable, as it relies on manipulating the loss computation through malicious code. Specifically, in a benign training pipeline, where the loss computation is not compromised, the blind backdoor cannot function as intended, even if the sourced pre-trained weights are backdoored. In contrast, our proposed latent backdoor remains effective by leveraging controlled manipulation of a single task during pre-training, embedding a dormant backdoor that activates during the benign training of subsequent tasks. Finally, we note that detecting changes in pre-trained weights is challenging due to the randomness introduced by factors like the optimizer, initial seed, and data variations. In Section 6, we present empirical results on how these attacks evade backdoor detection mechanisms.
@@
In summary, for BTB, we assume that the adversary compromises the loss computation in the ML code. Although the adversary may have knowledge of the architecture or data domain, they lack access to the training data and have no knowledge of the hyperparameters. In contrast, for LTB, the adversary controls the training process of a single task.
@@ 5.1 Blind Task Backdoor
As mentioned in Section 4.1, in Blind Task Backdoor (i.e., blind backdoor), the adversary generates the attack code by manipulating the loss value computation of the ML code, indirectly impacting the training process of the model across all tasks. For this attack to be successful, the modification of the loss computation should achieve a high attack success rate on the triggered inputs on the current and future tasks, and maintain high classification accuracy on the clean inputs across all tasks. To achieve this, we devise a customized blind loss, aiming to effectively solve this multi-objective optimization problem, as detailed in Eqn. 1:
@@ 5  Compute $F\leftarrow\nabla_{\theta}^{2}\ell$ for each $\theta$
6 $\triangleright$ Select Stable Neurons from $n$ Neurons
@@ 11  $\ell_{latent}\leftarrow L_{i}(\theta(x^{+}),y^{+})+\text{ReLU}(L_{i}(\theta(x),y)-\epsilon)$
$\underset{\theta,\lambda,\mu}{\min}\,(\ell_{blind}^{i}+\sum_{j}\lambda_{j}\delta+\frac{\mu}{2}\sum_{j}\delta^{2}),$ (3)
@@ 5.2 Latent Task Backdoor
DFM analysis enables us to evaluate neurons’ contributions to the classification performance of the task controlled by the adversary by assessing the loss function’s sensitivity to variations in neurons. Interestingly, our in-depth analysis revealed a significant overlap between the stable neurons and those identified by DFM analysis as crucial for the task’s classification performance. As a result, we select the top important neurons based on the importance matrix $F$, forming the candidate set, with $\kappa$ approximately corresponding to the 98th percentile of $F$ in our experiments. To maintain the attack’s stealth and persistence, we select a proportion $P\%$ of the candidate set for trigger embedding (line 5). Our algorithmic-level observations in Section 3 revealed that replay-based algorithms experience more significant parameter changes, necessitating a higher percentage of neurons for trigger embedding in replay-based algorithms. In our experiments, we set $P$ to 70% and 90% of the candidate set for regularization-based and replay-based algorithms, respectively.
@@ 6 Experiments
To validate the effectiveness of our proposed blind backdoor (BTB) and latent backdoor (LTB) attacks, we conducted extensive experiments (details are in Table 4 in the Appendix) to evaluate: (i) static backdoor against six CL algorithms, (ii) dynamic backdoor (iii) physical backdoor, (iv) backdoor on text classification, (v) comparative analysis against other backdoor methods, (vi) backdooring different tasks in CL, and (vii) evading existing backdoor detection defenses [^1] Code is available on https://doi.org/10.5281/zenodo.14728872. We perform backdoor attacks on task-based CL with the code developed based on [47]. The backdoor is embedded in task 1 unless stated otherwise. All experiments are conducted using PyTorch and executed on an NVIDIA A100 GPU from ACCESS CI [6]. For BTB, we set the number of iterations ($n$) to 300, $\alpha$ to 0.0010.0010.001, $\beta$ to 0.00010.00010.0001, and $\tau$ to be $0.05\times\ell_{j}$, where $\ell_{j}$ is the loss value of original task $j$. For the LTB attack, we set $v^{tigger}$ to 0.5, $\epsilon$ to 0.1 of the backdoor loss ($0.1\times L(\theta(x^{+}),y^{+})$), $\kappa$ to 98 percentile of all neurons, and $P$ to 70% and 90% for regularization-based and replay-based algorithms.
@@
CNN SplitMNIST 0.950.950.95 ±0.05$\pm$ 0.05± 0.05 0.920.920.92 ±0.02$\pm$ 0.02± 0.02 0.930.930.93 ±0.03$\pm$ 0.03± 0.03 0.910.910.91 ±0.02$\pm$ 0.02± 0.02 0.940.940.94 ±0.01$\pm$ 0.01± 0.01 0.910.910.91 ±0.03$\pm$ 0.03± 0.03 0.910.910.91 ±0.05$\pm$ 0.05± 0.05 0.920.920.92 ±0.02$\pm$ 0.02± 0.02 0.920.920.92 ±0.04$\pm$ 0.04± 0.04 0.930.930.93 ±0.03$\pm$ 0.03± 0.03 0.910.910.91 ±0.07$\pm$ 0.07± 0.07 0.930.930.93 ±0.03$\pm$ 0.03± 0.03
PermutedMNIST 0.930.930.93 ±0.04$\pm$ 0.04± 0.04 0.820.820.82 ±0.09$\pm$ 0.09± 0.09 0.850.850.85 ±0.09$\pm$ 0.09± 0.09 0.800.800.80 ±0.05$\pm$ 0.05± 0.05 0.900.900.90 ±0.03$\pm$ 0.03± 0.03 0.820.820.82 ±0.01$\pm$ 0.01± 0.01 0.920.920.92 ±0.06$\pm$ 0.06± 0.06 0.850.850.85 ±0.03$\pm$ 0.03± 0.03 0.910.910.91 ±0.04$\pm$ 0.04± 0.04 0.26 ±0.02$\pm$ 0.02± 0.02 0.890.890.89 ±0.03$\pm$ 0.03± 0.03 0.810.810.81 ±0.04$\pm$ 0.04± 0.04
SplitCIFAR10 0.910.910.91 ±0.02$\pm$ 0.02± 0.02 0.760.760.76 ±0.01$\pm$ 0.01± 0.01 0.910.910.91 ±0.03$\pm$ 0.03± 0.03 0.830.830.83 ±0.04$\pm$ 0.04± 0.04 0.880.$8$0.04$\pm$ 0.04± 0.04 0.850.850.85 ±0.01$\pm$ 0.01± 0.01 0.840.840.84 ±0.04$\pm$ 0.04± 0.04 0.810.810.81 ±0.03$\pm$ 0.03± 0.03 0.940.940.94 ±0.03$\pm$ 0.03± 0.03 0.52 ±0.02$\pm$ 0.02± 0.02 0.890.890.89 ±0.06$\pm$ 0.06± 0.06 0.860.860.86 ±0.03$\pm$ 0.03± 0.03
ResNet18 SplitMNIST 0.930.930.93 ±0.01$\pm$ 0.01± 0.01 0.960.960.96 ±0.03$\pm$ 0.03± 0.03 0.910.910.91 ±0.01$\pm$ 0.01± 0.01 0.920.920.92 ±0.04$\pm$ 0.04± 0.04 0.930.930.93 ±0.03$\pm$ 0.03± 0.03 0.910.910.91 ±0.02$\pm$ 0.02± 0.02 0.910.910.91 ±0.01$\pm$ 0.01± 0.01 0.980.980.98 ±0.04$\pm$ 0.04± 0.04 0.890.890.89 ±0.03$\pm$ 0.03± 0.03 0.910.910.91 ±0.02$\pm$ 0.02± 0.02 0.900.900.90 ±0.03$\pm$ 0.03± 0.03 0.930.930.93 ±0.04$\pm$ 0.04± 0.04
PermutedMNIST 0.930.930.93 ±0.01$\pm$ 0.01± 0.01 0.860.860.86 ±0.06$\pm$ 0.06± 0.06 0.890.890.89 ±0.03$\pm$ 0.03± 0.03 0.880.$8$0.04$\pm$ 0.04± 0.04 0.920.920.92 ±0.01$\pm$ 0.01± 0.01 0.790.790.79 ±0.03$\pm$ 0.03± 0.03 0.930.930.93 ±0.06$\pm$ 0.06± 0.06 0.910.910.91 ±0.07$\pm$ 0.07± 0.07 0.910.910.91 ±0.01$\pm$ 0.01± 0.01 0.35 ±0.06$\pm$ 0.06± 0.06 0.850.850.85 ±0.03$\pm$ 0.03± 0.03 0.850.850.85 ±0.01$\pm$ 0.01± 0.01
SplitCIFAR10 0.900.900.90 ±0.03$\pm$ 0.03± 0.03 0.800.800.80 ±0.04$\pm$ 0.04± 0.04 0.900.900.90 ±0.03$\pm$ 0.03± 0.03 0.820.820.82 ±0.02$\pm$ 0.02± 0.02 0.820.820.82 ±0.02$\pm$ 0.02± 0.02 0.780.780.78 ±0.03$\pm$ 0.03± 0.03 0.900.900.90 ±0.01$\pm$ 0.01± 0.01 0.830.830.83 ±0.03$\pm$ 0.03± 0.03 0.890.890.89 ±0.02$\pm$ 0.02± 0.02 0.63 ±0.03$\pm$ 0.03± 0.03 0.930.930.93 ±0.01$\pm$ 0.01± 0.01 0.880.$8$0.07$\pm$ 0.07± 0.07